What are the possible values of the third side of a triangle if the measure of two of its sides are 9 and 14?
Answer by checkley77(12569)

9^2+14^2=x^2
81+196=x^2
x^2=277
x=sqrt277
x=16.6433 answer.
9^2+x^2=14^2
81+x^2=196
x^2=196-81
x^2=115
x=sqrt115
x=10.7238 answer.